# DiLi: A Lock-Free Asynchronously Distributable Linked List

### v1

@@ 9 Concluding Remarks
Although designed as a linked list, the implementation of DiLi provides a performance comparable to skip lists [11, 28, 10] on a single machine. Additionally, while these skip list implementations cannot provide horizontal scaling via distribution, DiLi is capable of providing such scaling. An implication of this is that any lock-free data structure that is based off of Harris List, can easily be converted into a high-performance distributed data structure using the techniques in this paper.
The data structure considered in this paper, a linked list, differs from other data structures such as hash tables in one important way. The elements in the hash table are not connected to each other. In contrast, elements in a linked list are connected to each other by the notion of the next element. These relations are preserved by the split and move operations in DiLi. It is therefore possible to ensure that these operations can be added to DiLi while ensuring that their lock-free nature is preserved even while the list is being split or moved. This will open up the possibility of horizontally scaling other lock-free data structures (e.g., trees) while preserving their lock-free nature.
@@ Appendix A Registry Operations
398
399Registry *registry;
400 void addEntry(Entry *entry):
@@
402  do
403     currReg = registry;
404     Registry newReg;
405     for $i=0$; $(i<currReg\rightarrow size\wedge$ $currReg\rightarrow entries[i]\rightarrow keyMin<newKey)$; $i=i+1$ do
@@
409     for $(;i<currReg\rightarrow size;i=i+1)$ do
@@
411
412    i = i + 1;
[…]
415
[…]
417Entry *getByKey(Key key):
418  currReg == registry;
419  $\text{entries}=\text{currReg}\rightarrow\text{entries}$;
420
[…]
422  while left $\leq$ right do
423     middle == (left ++ right)//2;
[…]
425        right == middle-1;
426
427    else
[…]
429           left == middle ++1;
430
431       else
432           return entries[middle];
433
434 return null;
435
436
Algorithm 6 Registry Operations
As depicted in Algorithm˜6, the Registry features two main operations – getByKey(key) and addEntry(entry). Add Entry copies the registry to a new registry, while placing the new entry at the right position of the array, that is kept sorted by the keyMin of the entries. In our implementation, we additionally use [25] to safely reclaim old registry (memory stored currReg pointer variable in the addEntry() routine) after the last pointer to it goes out of scope. The operation getByKey() reads the value of the current registry pointer and does the classic binary search on the registry as an array. Note that since we allow multi read (used by client operations), single write (used only by Split) for the registry, any implementation of a lock-based binary search trees that has a wait free search could be used instead of our simple illustration with arrays and copy-on-writes. A removeEntry() method can be implemented by a similar copy-on-write technique.
Appendix B Merge Operation
The algorithm for the Merge operation is depicted in Algorithm˜7. The operation takes the two registry entries of the sublists that are in the same server and next to one another as input, then merges the right sublist with the left sublist and returns the left sublist entry as the response. The algorithm is much similar to Split, but the logic happens in a different order. First, The left sublist entry is updated with the merged key range. Then the right sublist entry is removed from the index. This allows subsequent traversals of the client operations to be possible just from the left sublist. Then the stCt and endCt of the items in the right sublist are updated with the left sublist counters.
Now all that is left before announcing the merge is to remove the subtail and subhead present at the middle of the merged sublist. This is done using a restricted double compare single swap (RDCSS) that is implemented using 3 CAS operations [15]. The operation takes two pairs of <variable, value> pairs (hence the first 4 parameters) for comparison and replaces the first variable with the fifth parameter. For the merge to remove the subtail subhead block, the operation repeatedly tries to update the next pointer of the last element of the left sublist to the first element of the right sublist. To make it asynchronous, it must succeed only when no insertions take place on the subhead to be removed. Hence, the double comparison is performed for a single swap. After this, the new offset is computed the same way as Split (Section˜5.3, by finding the difference between the counters during a momentary write-free period. The updated index is then announced to the other machines via RegisterMergedSublist message.
437Entry *Merge(Entry * leftEntry, Entry rightEntry):
438
@@ Appendix B Merge Operation
442
443 registry.removeEntry(rightEntry);
444
@@
446  do
447     prev = curr;
@@
460
[…]
462 do
463     $a_{1}$ = $\rightarrowstCt- leftEntry\rightarrowendCt$ ;
464     $a_{2}$ = $\rightarrowstCt- rightEntry\rightarrowendCt$ ;
465
@@
471
472 return leftEntry;
473
474bool RegisterMergedSublistRecv(Key keyMid, Ref SH):
475  Entry rightEntry = registry.getByKey(keyMid);
476  Entry leftEntry = registry.getByKey(keyMid - 1);
@@
478  registry.removeEntry(rightEntry);
479  return true;
480
@@ Appendix C Supplementary Results
Figure 4: Scatter plot of the time taken by background operations, plotted against the local elapsed time at which the opertion completed.
Figure˜4 shows the time taken by Split and Move operations (Move includes Switch) at the server that initially owned the entire key range for the experiment. The experiment was a workload of insert operations, that loaded 1M keys into two 2-core machines, with one of them not owning any part of the list key range at the beginning, and begins to receive sublists to own through Move operations the first minute of the load test. We observe that the load of insertions completed at the 180 second mark, after which no Split operations were performed (since all sublists are now within the threshold limit of 125 and the list is no longer receiving updates). After this time, only Moves were performed to complete load balancing.
@@ Appendix D Proof of Correctness for the Background Operation Algorithms
D.1 Properties of stCt and endCt of a Sublist for Delegations
@@
Proof D.2.
For any sublist, a subhead pointer that stores an item with stCt as negative, would redirect a request either using their newLoc or a fresh registry lookup. However, during a Move, there will exist two subheads with stCt being non-negative – 1. The subhead pointer (SH) of the source machine sublist copy. 2. The subhead pointer (SH*) created in the new machine sublist copy. No registry entry stores SH* during the Move phase. Hence every request gets routed through SH, and will continue to treat it as the active subhead since its stCt is not negative. During a Switch, SH will have its stCt as negative, and thus reroute all requests from the outdated registry entry to the newLoc, which is SH*. Registry entry is then updated to SH*, after which rerouting via SH will no longer be required and SH* will be the only active subhead.
@@
Proof D.4.
Wen stCt is incremented without incrementing endCt, the difference between the two is more than the offset. This makes it impossible for the computed CAS value check (line 4) to be true until the endCt is also incremented. When the stCt is already set to -$\infty$, no amount of increment can make it positive, thereby making the sublist copy inactive to perform.
D.2 Bound on the Number of Delegations
@@
During a Switch operation, a request to the old machine gets delegated to the new machine, bringing about an additional network round trip to the computation. Such a request has to be served in the new machine, as the distributed sublist assumes that another Switch begins long after all existing requests to the old machine for the sublist are completed. Thus, the maximum number of network round trips of a client operation during a Switch is atmost 3.
@@
D.3 Correctness of the Replay Algorithm in Move
Section˜5.4 shares the intuition behind the replay. The replay algorithm to recreate the moving sublist is done by reimagining the sublist to be built by InsertAfter (prevItem, newItem), instead of Insert(key), as it offers the exact item at which the insertion took place. Note that prevItem and newItem have sId, ts, key members present inside them, for easy notations in the proof. This distinguishes multiple items with the same key that were entered and removed, and ensures that if the prevItem had the same key as an old deleted item, then the replay waits till the newItem is inserted using the logical time stamp. Similarly, Remove(key) is replaced by Delete (prevItem).
@@ D.3 Correctness of the Replay Algorithm in Move
Proof D.8.
The lower timestamp clause of the Lemma follows from the fact that the insertion that lost the race in CAS will re-increment the logical clock to get a timestamp higher than the previous values. The latter part of the Lemma follows from the fact that the item from the latter $\tt{InsertAfter}$} will occupy the next pointer of A and will have its own next pointer pointing to the element that was inserted at A earlier.
@@
Proof D.10.
This follows from the fact that A must have existed in the sublist before $\tt{InsertAfter}$} (Ref(A), C) was called.
@@
Proof D.12.
This follows from the fact that the entire linked list structure (excluding next pointer mark values) has been created by the use of $\tt{InsertAfter}$} and $\tt Split$} operations. Since $\tt Split$} only inserts special, client-invisible nodes($\tt SubHead$} and $\tt SubTail$}), it is safe to argue that the nodes A and B must have been inserted into the list through $\tt{InsertAfter}$} operations. $\tt{InsertAfter}$} can only insert nodes to the right side of the node provided to it. Hence, if node B is appearing at the right of A, then either it was inserted at A or was inserted at an ancestor of A, at a time before A was inserted into the list. Thus, in the former case, $B.ts>A.ts$ and in the latter case, $B.ts<A.ts$.
@@
Proof D.14.
This follows from Lemma˜D.11, by which B was not inserted at A in this case. Hence B is not a competing insertion at A, and C can be inserted just after A as the first replicated insertion at A.
@@
Proof D.16.
From Lemma˜D.11, this conveys that B was inserted at A. This means C and B were possibly competing to insert at A, and so if $C.ts<B.ts$, C has to be present after B, according to Lemma˜D.7. After B, there could exist other descendants of A that were inserted in ways similar to B. Hence, the replay has to find the first node D after A, such that $C.ts>D.ts$, which will ensure either that D is not a descendant of A or that D is a descendant of A that was inserted into the list before C was inserted. Either case, C has to be inserted just before D.
@@
Proof D.18.
Combining Lemma˜D.13 and Lemma˜D.15, the algorithm to replay a request of RepInsert(prevItem,item) condenses down to finding the first node(currcurr in Line 4 of Algorithm˜4), such that $curr.ts<item.ts$ (as in Line 4). This will ensure that currcurr is either a node of type B in Lemma˜D.13 or a node of type D in Lemma˜D.15. Hence, the insert replays reconstruct the same sublist that the corresponding $\tt{InsertAfter}$} operations had created on the other server.
D.4 Termination Condition of the Asynchronous Background Operations
While we will be showcasing the practical termination of the background operations in Section˜7, and proved its asynchronous nature on client operations through lock-freedom, we still identify their termination conditions in this subsection.
Split: While the correctness of Split is straightforward from the explanation in Section˜5.3, the termination of Split operation depends on overcoming three different race conditions – 1. It first competes with the insert operations (Line 3 on $sItem$ to insert the “ST-SH” block into the list. 2. After insertion, it loops to find the new offsets for the split sublists (Line 3). This requires a brief moment in time when no updates take place on either sublist. 3. Finally, it competes with other background Split operations on other sublists to update the registry (Lines 3 to 3).
Move: The correctness of Move comes from the exhaustive proof of correctness of the Replay Algorithm (provided in Section˜D.3) to generate a live clone of the sublist. However, its termination depends on only one race condition - setting stCt to -$\infty$ through the CAS in Line 4 (succeeding this CAS initiates the Switch). This is a type of spin lock, that requires only a brief moment when there are no ongoing updates to the sublist.
Switch: Correctness of Switch comes from the single active subhead pointer property (as in Section˜D.1). However, it however has one race condition – It needs to find the latest subtail of the previous sublist (if one exists) and change its next pointer (Line 5 to 5). This thus competes with Move and Switch operations of that previous sublist, which is practically trivial to overcome as Move of the same sublist should anyway not be invoked too often to avoid performance loss from too many delegations of client operations on this sublist.

### v2

@@ 9. Concluding Remarks
Although designed as a linked list, the implementation of DiLi provides a performance comparable to skip lists (Fraser, 2004; Sundell and Tsigas, 2004; Fomitchev and Ruppert, 2004a) on a single machine and even surpasses them as a library by up to 26% throughput in write-heavy workloads. While these skip list implementations cannot provide horizontal scaling via distribution, DiLi is additionally capable of providing such scaling. Thus, DiLi provides a foundation for converting other high performance single machine data structures into distributed data structures. While the experiments illustrates a simple usage of the $\tt Split$} and $\tt Move$} primitives for YCSB workloads, a future work can be designing strategies for any generic workload (such as TPC and synchrobench benchmarks.
The data structure considered in this paper, a linked list, differs from other data structures such as hash tables in one important way. The elements in the hash table are not connected to each other. In contrast, elements in a linked list are connected to each other by the notion of the next element. These relations are preserved by the split and move operations in DiLi. It is therefore possible to ensure that these operations can be added to DiLi while ensuring that their lock-free nature is preserved even while the list is being split or moved. This will open up the possibility of horizontally scaling other lock-free data structures (e.g., trees) while preserving their lock-free nature. Another future direction would be to introduce operations that span multiple machines and exploit the linked list structure - such as ranged queries and snapshots. We believe this paper will be foundational for such future work.
Finally, this paper focused on the scalability aspect of distribution while maintaining lock-freedom of client operations. Another important direction would be fault tolerance, such as additional maintenance of the list data in a replication fashion.
@@ Appendix A Registry Operations
Registry *registry;
void addEntry(Entry *entry):
@@
  do
     currReg = registry;
     Registry newReg;
     for $i=0$; $(i<currReg\rightarrow size\wedge$ $currReg\rightarrow entries[i]\rightarrow keyMin<newKey)$; $i=i+1$ do
@@
     for $(;i<currReg\rightarrow size;i=i+1)$ do
@@
     end for
    i = i + 1;
@@
end
Entry *getByKey(Key key):
  currReg == registry;
  $\text{entries}=\text{currReg}\rightarrow\text{entries}$;
@@
  while left $\leq$ right do
     middle == (left ++ right)//2;
[…]
        right == middle-1;
    else
[…]
           left == middle ++1;
       else
           return entries[middle];
        end if
     end if
  end while
 return null;
end
Algorithm 3 Registry Operations
As depicted in Algorithm 3, the Registry features two main operations – getByKey(key) and addEntry(entry). Add Entry copies the registry to a new registry, while placing the new entry at the right position of the array, that is kept sorted by the keyMin of the entries. In our implementation, we additionally use (Michael, 2004) to safely reclaim old registry (memory stored currReg pointer variable in the addEntry() routine) after the last pointer to it goes out of scope. The operation getByKey() reads the value of the current registry pointer and does the classic binary search on the registry as an array. Note that since we allow multi read (used by client operations), single write (used only by Split) for the registry, any implementation of a lock-based binary search trees that has a wait free search could be used instead of our simple illustration with arrays and copy-on-writes. A removeEntry() method can be implemented by a similar copy-on-write technique.
Appendix B Merge Operation
The algorithm for the Merge operation is depicted in Algorithm 4. The operation takes the two registry entries of the sublists that are in the same server and next to one another as input, then merges the right sublist with the left sublist and returns the left sublist entry as the response. The algorithm is much similar to Split, but the logic happens in a different order. First, The left sublist entry is updated with the merged key range. Then the right sublist entry is removed from the index. This allows subsequent traversals of the client operations to be possible just from the left sublist. Then the stCt and endCt of the items in the right sublist are updated with the left sublist counters.
Now all that is left before announcing the merge is to remove the subtail and subhead present at the middle of the merged sublist. This is done using a restricted double compare single swap (RDCSS) that is implemented using 3 CAS operations (Harris et al., 2002). The operation takes two pairs of ¡variable, value¿ pairs (hence the first 4 parameters) for comparison and replaces the first variable with the fifth parameter. For the merge to remove the subtail subhead block, the operation repeatedly tries to update the next pointer of the last element of the left sublist to the first element of the right sublist. To make it asynchronous, it must succeed only when no insertions take place on the subhead to be removed. Hence, the double comparison is performed for a single swap. After this, the new offset is computed the same way as Split (Section 5.3, by finding the difference between the counters during a momentary write-free period. The updated index is then announced to the other machines via RegisterMergedSublist message.
Entry *Merge(Entry * leftEntry, Entry rightEntry):
@@ Appendix B Merge Operation
        $leftLast\rightarrownext$;
     end while
@@
 do
     $a_{1}$ = $\rightarrowstCt- leftEntry\rightarrowendCt$ ;
     $a_{2}$ = $\rightarrowstCt- rightEntry\rightarrowendCt$ ;
@@
[…]
 for $i\in(serverList-\{me\})$ do
[…]
  end for
 return leftEntry;
end
bool RegisterMergedSublistRecv(Key keyMid, Ref SH):
  Entry rightEntry = registry.getByKey(keyMid);
  Entry leftEntry = registry.getByKey(keyMid - 1);
[…]
  registry.removeEntry(rightEntry);
  return true;
end
@@ Appendix C Supplementary Results
C.1. Thread-Wise Variation of Various Write Workloads
Figure 4 shows the thread-wise results for all write workloads that we have experimented.
(a) Maximum average throughput achieved when write proportion is 10%.
(b) Maximum average throughput achieved when write proportion is 25%.
(c) Maximum average throughput achieved when write proportion is 50%.
(d) Maximum average throughput achieved when write proportion is 75%.
(e) Maximum average throughput achieved when write proportion is 90%.
Figure 4. Performance of DiLi and Skip List when used as a library to execute workloads of various write proportions. Key size is maintained to be 1 million throughout.
C.2. Thread-Wise Variation for Different Sizes of List
Figure 5 shows the thread-wise results for all sizes that we have experimented.
(a) Maximum size throughput achieved when size of the list is 100k.
(b) Maximum size throughput achieved when size of the list is 1m.
(c) Maximum size throughput achieved when size of the list is 2m.
(d) Maximum average throughput achieved when write proportion is 4m.
(e) Maximum size throughput achieved when size of the list is 5m.
Figure 5. Performance of DiLi and Skip List when used as a library to execute workloads of various key sizes. Write Proportion is maintained to be 50% throughout.
C.3. Practical Termination of Background Operations
Figure 6. Scatter plot of the time taken by background operations, plotted against the local elapsed time at which the opertion completed.
Figure 6 shows the time taken by Split and Move operations (Move includes Switch) at the server that initially owned the entire key range for the experiment. The experiment was a workload of insert operations, that loaded 1M keys into two 2-core machines, with one of them not owning any part of the list key range at the beginning, and begins to receive sublists to own through Move operations the first minute of the load test. We observe that the load of insertions completed at the 180 second mark, after which no Split operations were performed (since all sublists are now within the threshold limit of 125 and the list is no longer receiving updates). After this time, only Moves were performed to complete load balancing.
@@ Appendix D Proof of Correctness for the Background Operation Algorithms
D.1. Properties of stCt and endCt of a Sublist for Delegations
@@
Proof.
For any sublist, a subhead pointer that stores an item with stCt as negative, would redirect a request either using their newLoc or a fresh registry lookup. However, during a Move, there will exist two subheads with stCt being non-negative – (1) The subhead pointer (SH) of the source machine sublist copy. (2) The subhead pointer (SH*) created in the new machine sublist copy. No registry entry stores SH* during the Move phase. Hence every request gets routed through SH, and will continue to treat it as the active subhead since its stCt is not negative. During a Switch, SH will have its stCt as negative, and thus reroute all requests from the outdated registry entry to the newLoc, which is SH*. Registry entry is then updated to SH*, after which rerouting via SH will no longer be required and SH* will be the only active subhead. ∎
@@
Proof.
Wen stCt is incremented without incrementing endCt, the difference between the two is more than the offset. This makes it impossible for the computed CAS value check (line 2) to be true until the endCt is also incremented. When the stCt is already set to -$\infty$, no amount of increment can make it positive, thereby making the sublist copy inactive to perform. ∎
D.2. Bound on the Number of Delegations
@@
During a Switch operation, a request to the old machine gets delegated to the new machine, bringing about an additional network round trip to the computation. Such a request has to be served in the new machine, as the distributed sublist assumes that another Switch begins long after all existing requests to the old machine for the sublist are completed. Thus, the maximum number of network round trips of a client operation during a Switch is atmost 3. ∎
@@
D.3. Correctness of the Replay Algorithm in Move
Section 5.4 shares the intuition behind the replay. The replay algorithm to recreate the moving sublist is done by reimagining the sublist to be built by InsertAfter (prevItem, newItem), instead of Insert(key), as it offers the exact item at which the insertion took place. Note that prevItem and newItem have sId, ts, key members present inside them, for easy notations in the proof. This distinguishes multiple items with the same key that were entered and removed, and ensures that if the prevItem had the same key as an old deleted item, then the replay waits till the newItem is inserted using the logical time stamp. Similarly, Remove(key) is replaced by Delete (prevItem).
@@ D.3. Correctness of the Replay Algorithm in Move
Proof.
The lower timestamp clause of the Lemma follows from the fact that the insertion that lost the race in CAS will re-increment the logical clock to get a timestamp higher than the previous values. The latter part of the Lemma follows from the fact that the item from the latter $\tt{InsertAfter}$} will occupy the next pointer of A and will have its own next pointer pointing to the element that was inserted at A earlier. ∎
@@
Proof.
This follows from the fact that A must have existed in the sublist before $\tt{InsertAfter}$} (Ref(A), C) was called. ∎
@@
Proof.
This follows from the fact that the entire linked list structure (excluding next pointer mark values) has been created by the use of $\tt{InsertAfter}$} and $\tt Split$} operations. Since $\tt Split$} only inserts special, client-invisible nodes($\tt SubHead$} and $\tt SubTail$}), it is safe to argue that the nodes A and B must have been inserted into the list through $\tt{InsertAfter}$} operations. $\tt{InsertAfter}$} can only insert nodes to the right side of the node provided to it. Hence, if node B is appearing at the right of A, then either it was inserted at A or was inserted at an ancestor of A, at a time before A was inserted into the list. Thus, in the former case, $B.ts>A.ts$ and in the latter case, $B.ts<A.ts$. ∎
@@
Proof.
This follows from Lemma D.6, by which B was not inserted at A in this case. Hence B is not a competing insertion at A, and C can be inserted just after A as the first replicated insertion at A. ∎
@@
Proof.
From Lemma D.6, this conveys that B was inserted at A. This means C and B were possibly competing to insert at A, and so if $C.ts<B.ts$, C has to be present after B, according to Lemma D.4. After B, there could exist other descendants of A that were inserted in ways similar to B. Hence, the replay has to find the first node D after A, such that $C.ts>D.ts$, which will ensure either that D is not a descendant of A or that D is a descendant of A that was inserted into the list before C was inserted. Either case, C has to be inserted just before D. ∎
@@
Proof.
Combining Lemma D.7 and Lemma D.8, the algorithm to replay a request of RepInsert(prevItem,item) condenses down to finding the first node(currcurr in Line 2 of Algorithm 2), such that $curr.ts<item.ts$ (as in Line 2). This will ensure that currcurr is either a node of type B in Lemma D.7 or a node of type D in Lemma D.8. Hence, the insert replays reconstruct the same sublist that the corresponding $\tt{InsertAfter}$} operations had created on the other server. ∎
D.4. Termination Condition of the Asynchronous Background Operations
While we will be showcasing the practical termination of the background operations in Section 7, and proved its asynchronous nature on client operations through lock-freedom, we still identify their termination conditions in this subsection.
Split: While the correctness of Split is straightforward from the explanation in Section 5.3, the termination of Split operation depends on overcoming three different race conditions – (1) It first competes with the insert operations (Line 1 on $sItem$ to insert the “ST-SH” block into the list. (2) After insertion, it loops to find the new offsets for the split sublists (Line 1). This requires a brief moment in time when no updates take place on either sublist. (3) Finally, it competes with other background Split operations on other sublists to update the registry (Lines 1 to 1).
Move: The correctness of Move comes from the exhaustive proof of correctness of the Replay Algorithm (provided in Section D.3) to generate a live clone of the sublist. However, its termination depends on only one race condition - setting stCt to -$\infty$ through the CAS in Line 2 (succeeding this CAS initiates the Switch). This is a type of spin lock, that requires only a brief moment when there are no ongoing updates to the sublist.
Switch: Correctness of Switch comes from the single active subhead pointer property (as in Section D.1). However, it however has one race condition – It needs to find the latest subtail of the previous sublist (if one exists) and change its next pointer (Line 2 to 2). This thus competes with Move and Switch operations of that previous sublist, which is practically trivial to overcome as Move of the same sublist should anyway not be invoked too often to avoid performance loss from too many delegations of client operations on this sublist.